The inverse demand function is given by P = 2 – 0.13 Q.

What is the price elasticity of demand?

What is the price elasticity of demand when P = 1.2


1
Expert's answer
2021-03-08T07:20:20-0500

We can find the price elasticity of demand as follows:


"\epsilon=-\dfrac{P}{Q}\dfrac{dQ}{dP},""Q=\dfrac{2-P}{0.13}, \dfrac{dQ}{dP}=-\dfrac{1}{0.13},""\epsilon=-\dfrac{0.13P}{2-P}\cdot-(\dfrac{1}{0.13})=\dfrac{P}{2-P}."

The price elasticity of demand when P = 1.2 equals:


"\epsilon=\dfrac{1.2}{2-1.2}=1.5"